OpenStudy (anonymous):

Mrs. Waight graphed the following system of linear equations: 3x + y= -1 2x - y = -4 Which of the following is a solution of the system of equations? A. (-1, 2) B. (1, -2) C. (2, -1)

OpenStudy (anonymous):

A is my choice.

OpenStudy (eric_d):

Do you know how to find value of x

geerky42 (geerky42):

A is correct.

OpenStudy (eric_d):

3x+y=-1 y=-1-3x 2x-y=-4 2x+4=y y=2x+4 -1-3x=2x+4 -1-4=2x+3x -5=5x x=-1 Yes, it is A